OpenStudy (anonymous):

Eliminate the parameter. x = 5t, y = t + 8 y= 5x + 8 y = x/5 + 8 y = 5x - 8 y = x/5 - 8 I choose the last one as my answer. Is that right?

OpenStudy (anonymous):

@mathmath333 can you help?

OpenStudy (anonymous):

yes thats right

OpenStudy (anonymous):

ya thats correct good job SunShine201

OpenStudy (anonymous):

Thank you. Do you think you could check one more?

OpenStudy (surry99):

go ahead and post it.

OpenStudy (anonymous):

ya ill check it to

OpenStudy (anonymous):

Eliminate the parameter. x = t2 + 3, y = t2 - 2 y = x + 5, x ≥ 1 y = x2 - 5, x ≥ 1 y = x - 5, x ≥ 1 y = x2 + 5, x ≥ 1

OpenStudy (anonymous):

I think it's C but I also think it could be A. Which is it?

OpenStudy (surry99):

How are you trying to solve it?

OpenStudy (surry99):

The method matters more than the answer.

OpenStudy (anonymous):

hey SunShine201 check out this link i posted it helps my friend realy realy good http://www.wolframalpha.com/input/?i=x+%3d+t%5e2+%2b+3%2c+y+%3d+t%5e2+-+2&incTime=true

OpenStudy (anonymous):

I don't know why the sign didn't show up but it is suppose to be  for each answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

anyway i got this since x = 2t^2 + t^2, then x = 3 t^2 So t = (1/3 x) ^0.5 So y = 2/3 x + (1/3 x) ^1.5

OpenStudy (anonymous):

hope this helps you

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

np brb